A firm faces the production function below and can buy L at Php240 a unit and K at Php4 a unit.

A. If it has a budget of Php 16,000 what combination of K and L should it use to maximize output?

Q = 2K 0.2 L 0.6


1
Expert's answer
2021-07-01T13:50:06-0400

"Q= 2K^{0.2}L^{0.6} ............ (1)"

Wage w = Php 240

rent = Php 4 

Budegt = 16000


Budegt line 

"wL+rK = budegt"  

Putting values 

"240L+4K = 16000 ....... (2)"

Slope of the isocost line "= -\frac{240}{4}"  

Marginal product of labour "MPL = 2(0.6)K^{0.2}L^{-0.4}"   

Marginal product of labour "MPK = 2(0.2)K^{-0.8}L^{0.6}"

Slope of the isoquants "=-MRTS =\frac {-MPL}{MPK}"


"MRTS =\frac {-2(0.6)K^{0.2}L^{-0.4}}{ 2(0.2)K^{-0.8}L^{0.6} }\\=-\frac{3K}{L}"


Condition of optimality 

The slope of isoquants = slope of the isocost line 

"-\frac{3K}{L}=-\frac{240}{4}\\\frac{3K}{L}=60\\K=20L"


Putting value of K in isocost equation 

"240L+4(20L) = 16000"

"i.e. 320L=16000"

"i.e. L =50"

So "K = 20(50) = 1000"

If K = 1000 and L =50, then we get maximum output.